Calculate the gains to the consumer from the subsidy if the demand and supply of Chinese coal are
Qd = 60 − 0.5P.
Qs = −2 + 2P.
1
Expert's answer
2020-11-12T10:35:15-0500

60-0.5p=-2+2p

2.5p=62

p=24.8

Q=-2+2*24.8

Q=47.6